# O0.45cmO0.1cm =

### v1

@@ Abstract
Effective versions of strong measure zero sets are developed for various levels of complexity and computability. It is shown that the sets can be equivalently defined using a generalization of supermartingales called odds supermartingales, success rates on supermartingales, predictors, and coverings. We show Borel’s conjecture of a set having strong measure zero if and only if it is countable holds in the time and space bounded setting. At the level of computability this does not hold. We show the computable level contains sequences at arbitrary levels of the hyperarithmetical hierarchy by proving a correspondence principle yielding a condition for the sets of computable strong measure zero to agree with the classical sets of strong measure zero.
@@ 1 Introduction
Measure theory is a well studied area of mathematics resulting in a notion of smallness for sets of measure zero. Resource bounded versions of measure zero sets were developed by Lutz [18]. Hausdorff [8] refined measure zero sets in the classical (non effective) setting with dimension. This yields the dimension zero sets which are a subclass of the measure zero sets. An effective version of dimension defined by Lutz [20] along with effective measure have lead to many connections and results in areas including complexity and randomness. For more of an overview see [19, 22]. In this paper, we continue the refinement by developing effective versions of a subclass of dimension zero sets called strong measure zero.
Borel [4] defined a set $X\subseteq\mathbb{R}$ to be strong measure zero (SMZ) if for any sequence $(\epsilon_{n})_{n\in\mathbb{N}}$ of positive reals there is a set $(I_{n})_{n\in\mathbb{N}}$ of intervals such that
@@
where $|I_{n}|$ is the length of $I_{n}$. It is easy to see that every countable set has strong measure zero. Borel conjectured that a set is strong measure zero if and only if it is countable. The work of Sierpiński [31] and Laver [14] proved that this is independent of ZFC. Strong measure zero is easily defined for subsets of Cantor space which we focus on in this paper. For a thorough background on strong measure zero sets see [2].
@@
In this work we investigate requiring the supermartingales themselves to be effective, leading instead to a stronger notion than strong measure zero set in the classical setting. These supermartingale objects are similar to those used to define effective measure and dimension, but we will show in section 3 and that there are equivalent characterizations using the success rate of standard supermartingales.
@@
A natural effective version of Borel’s conjecture with time and space bounded resources along with at the computable level exists. This uses an effective countability definition by Lutz [18] which are rroughly sets of sequences than can be uniformly enumerated within the resource bound. We prove that the effectively countable sets have effective strong measure zero at every level. For the time and space bounded setting it is proven that Borel’s conjecture holds. However, it does not hold at the computable level.
We investigate algorithmic strong measure zero occurring at the level of lower-semicomputability. The main result is that the resulting class is equivalent to the well studied set of NCR reals defined by Reimann and Slaman [27] containing sequences that are not algorithmically random with respect to any continuous probability measure. We will discuss this connection more in section 5, but note for now that in the classical setting these are not equivalent. NCR has been proven to be countable [28]. We thus have the following implication diagram where $\mathrm{\Delta}_{R}$ represents a time or space bounded resource class and the dashed arrow implies independence from ZFC.
@@
We show that effective strong measure zero sets can be defined using effective coverings resulting in a characterization similar to Borel’s original definition (1.1). Using this, we prove a correspondence principle giving sufficient conditions for computable, algorithmic, and classical versions of strong measure zero to agree on a set. This result is analogous to Hitchcock’s [12] correspondence principle for effective dimension. We use our result to show that there are sequences at arbitrarily high levels of the hyperarithmetic hierarchy with computable strong measure zero, mirroring a result for NCR.
Many of the characterizations of strong measure zero sets involve success with respect to a limit superior. We therefore investigate what happens if success is required in the limit inferior. For dimension, Lutz [20] showed a limit superior condition on martingales characterizes Hausdorff dimension while Althreya, Hitchcock, Lutz and Mayordomo [1] showed packing dimension can be characterized with a limit inferior. Requiring limit inferior success at all gauges leads to sets with strong packing dimension zero (SPDZ). In the classical setting, the existence of uncountable strong packing dimension zero sets is also independent of ZFC, see for example [33]. For time and space restrictions the sets remain the same. However at the computable level, the sets now correspond with a weak computably-countable notion that we define. It is also shown that the algorithmic version contains exactly the decidable sequences, giving the following implication chart.
@@ 2 Preliminaries
We work in the Cantor space $\mathbf{C}=\{0,1\}^{\infty}$ consisting of all infinite binary sequences. A string is a finite, binary string $w\in\{0,1\}^{*}$ which has length $|w|$. We write $\lambda$ for the empty string of length 0. For $10\leq i\leq j\leq|w|-$ 1 we write $w[i\dots j]$ for the string consisting of the $i^{\text{th}}$ through $j^{\text{th}}$ bits of $w$. Similarly for $i,j\in\mathbb{N}$ with $i\leq j$, we let $S[i\dots j]$ be the string consisting of the $i^{\text{th}}$ through $j^{\text{th}}$ bits of $S$.
@@
[…]
@@
    $T=\bigcup_{n\in\mathbb{N}}\{f:\{0,1\}^{*}\to\{0,1\}^{*}|f\in\mathrm{DTIME}(g_{n})\}$.
Space bounded resource classes are defined analogously with DSPACE. In this paper we will let $\mathrm{\Gamma}_{R}$ be the set consisting of all time or space bounded resource classes using at least polynomial time or polynomial space.
@@
$\mathrm{all}$ =$=\{f:\{0,1\}^{*}\to\{0,1\}^{*}\}$
$\mathrm{comp}$ =$=\{f\in\mathrm{all}|f\text{ is computable}\}$
@@
The result of a constructor $\delta$ is the sequence ${R(\delta)\in\mathbf{C}}$ such that $\delta^{i}(\lambda)\sqsubseteq R(\delta)$ for all $i\in\mathbb{N}$.
@@
$R(\mathrm{\Delta})=\{R(\delta)|\delta\in\mathrm{\Delta}\text{ is a constructor}\}.$
@@
Similarly, for $\mathrm{\Delta}\in\mathrm{\Gamma}_{RCA}$ and a discrete domain $D$ we say that $f:D\to[0,\infty)$ is $\mathrm{\Delta}$-computable if there is an function $g:\mathbb{N}\times D\to\mathbb{Q}\cap[0,\infty)\in\mathrm{\Delta}$ called a computation of $f$ such that ${|g(r,x)-f(x)|\leq 2^{-r}}$ for all $r\in\mathbb{N}$. In the paper $\mathrm{\Gamma}$ denotes $\mathrm{\Gamma}_{RCA}$ together with the lower semicomputable class. We will also refer to the lower semicomputable level as algorithmic.
@@
where $F^{\prime}(f)=g$ is a computation of $F(f)$. For $\mathrm{\Delta}\in\mathrm{\Gamma}$ equal to comp or lower semicomputable, we say that the functional $F$ is in $\mathrm{\Delta}$ if there is an oracle Turing machine $M$ such that for all $f:A\to B,r\in\mathbb{N}$ and $x\in D$, $M$ on input $(r,x)$ with oracle access to $f$ outputs $g(r,x)$ where $g$ is a computation of $F(f)$ as defined above.
For time bounded $\mathrm{\Delta}$ we require that the time $M$ takes to output $g(x,r)$ is $h(|x|+r+\bar{g}(h(|x|+r))$ for some $h$ in $\mathrm{\Delta}$ where $\bar{g}(n)$ is the maximum length of the oracle $g$’s output on a string of length at most $n$. For example if $\mathrm{\Delta}$ is polynomial time then $M$ can query polynomially far out in the input length and use the maximum length of the queries as a parameter for the polynomial running time. Different notions of functional complexity exist, but all the results in the paper hold for any natural definition and most oracles will not impact the complexity in our use cases.
@@ 3.1 Strong Measure Zero
An odds function is any function $O:\{0,1\}^{*}\rightarrow[1,\infty)$. $O$ is said to be acceptable if $\prod_{w\sqsubset S}O(w)=\infty$ for all $S\in\mathbf{C}$.
@@
A set $X\subseteq\mathbf{C}$ has strong measure zero if and only if, for any acceptable odds $O:\{0,1\}^{*}\to[1,\infty)$, there is an $O$-supermartingale that succeeds on $X$.
@@
$\mu^{*}(X)=\inf\{\sum_{w\in A}\mu(w)|A\subseteq\{0,1\}^{*}\text{ and }X\subseteq\bigcup_{w\in A}C_{w}\}.$ (3.2)
@@
One can then show that there is an $O$-supermartingale that succeeds on a set $X$ if and only if $\mu_{O}^{*}(X)=0$. In their paper, they defined a set as having effective strong measure zero if for any computable odds function $O$ there exists an $O$-supermartingale that succeeds on the set. Here we instead use the following definition forcing effectivity on the $O$-supermartingales themselves.
@@
The fact that we limit odds functions to rationals in the interval $[1,2]$ may appear to be a weaker condition, but we will see that the definition is robust to this change and others in the next section. Note that all our odds functions therefore correspond to outer premeasures as in Lemma 3.3. From now on we will assume all odds functions are acceptable unless stated otherwise. We can also now see the connection between effective strong measure zero, measure zero, and dimension zero.
@@ 3.2 Strong Dimension Zero
$\limsup_{n\to\infty}\frac{d(S[0\dots n-1])}{2^{n}h(n)}=\infty$
@@
Suppose $X$ has $\mathrm{\Delta}$-strong measure zero with witnessing odds functional $F$. We will create a gauge functional $G$ as follows. Given a gauge function $h$, without loss of generality suppose $h(n)\leq 2h(n+1)$ for all $n$, otherwise $h$ can be transformed within $\mathrm{\Delta}$’s restrictions into an $h^{\prime}$ satisfying this with $h^{\prime}(n)\leq h(n)$ for all $n$. Now let $G$ use $F$ with odds function $O$ where $O(w)=\frac{h(|w|-1)}{h(|w|)}$ for all strings $w$ with $|w|>0$ and $h(-1)=1$. Then we have $\mu_{O}(w)=h(|w|)$ and
$G^{h}(w)=F^{O}(w)2^{|w|}h(|w|)$
@@
$\limsup_{n\to\infty}\frac{G^{h}(S[0\dots n])}{2^{n}h(n)}=\limsup_{n\to\infty}F^{O}(S[0\dots n])=\infty$
@@
$h(n)=\max\{\mu_{O}(w)|w\in\{0,1\}^{n}\}.$
Note that by compactness and the fact that $O$ is acceptable we have $lim_{n\to\infty}h(n)=0$. Therefore,
$F^{O}(w)=\frac{G^{h}(w)}{2^{|w|}\mu_{O}(w)}$
@@
$\limsup_{n\to\infty}F^{O}(S[0\dots n])\geq\limsup_{n\to\infty}\frac{G^{h}(S[0\dots n])}{2^{n}h(n)}=\infty$
@@
Let $D$ be any of the following sets or its intersection with $\mathbb{Q}$. Then, defining $\mathrm{\Delta}$ strong measure zero as the sets succeeding on all acceptable odds functions $O:\{0,1\}^{*}\to D$ results in equivalent definitions for any $\mathrm{\Delta}\in\mathrm{\Gamma}$.
@@ 3.3 Strong Predictability
$\liminf_{n\to\infty}\frac{L^{\log}_{\pi}(S[0\dots n-1])}{h(n)}<1$
@@
$\iff$ $\liminf_{n\to\infty}n+\log g(n)-\log d(S[0\dots n])=-\infty$
$\iff$ $\liminf_{n\to\infty}\mathcal{L}^{\log}_{\pi_{d}}(S[0\dots n])+\log g(n)=-\infty$
$\iff$ $\liminf_{n\to\infty}\frac{\mathcal{L}^{\log}_{\pi_{d}}(S[0\dots n])}{-\log g(n)}<1$
noting that the last equivalence is true since $\log g(n)$ is nonincreasing and goes to $\infty$- ∞ while $\mathcal{L}^{\log}_{\pi_{d}}(S[0\dots n])$ is positive. ∎
@@
Let $\mathrm{\Delta}\in\mathrm{\Gamma}$. A set $X\subseteq\mathbf{C}$ has $\Delta$-Hausdorff measure zero with respect to a gauge $g$ if and only if there is a $\mathrm{\Delta}$-computable superpredictor that $-\log(g)$-succeeds on $X$.

### v2

# Effective Versions of Strong Measure Zero

@@ Abstract
Effective versions of strong measure zero sets are developed for various levels of complexity and computability. It is shown that the sets can be equivalently defined using a generalization of supermartingales called odds supermartingales, success rates on supermartingales, predictors, and coverings. We show Borel’s conjecture that a set has strong measure zero if and only if it is countable holds in the time and space bounded setting. At the level of computability this does not hold. We show the computable level contains sequences at arbitrary levels of the hyperarithmetical hierarchy. This is done by proving a correspondence principle yielding a condition for the sets of computable strong measure zero to agree with the classical sets of strong measure zero.
@@ 1 Introduction
Measure theory is a well studied area of mathematics resulting in a notion of smallness for sets of measure zero. Resource bounded versions of measure zero sets were developed by Lutz [20]. Hausdorff [9] refined measure zero sets in the classical (non effective) setting with dimension. This yields the dimension zero sets which are a subclass of the measure zero sets. An effective version of dimension defined by Lutz [24] along with effective measure have led to many connections and results in areas including complexity and randomness. For more of an overview see [23, 25]. In this paper, we continue the refinement by developing effective versions of a subclass of dimension zero sets called strong measure zero.
Borel [4] defined a set $X\subseteq\mathbb{R}$ to be strong measure zero (SMZ) if for every sequence $(\epsilon_{n})_{n\in\mathbb{N}}$ of positive reals there is a set $(I_{n})_{n\in\mathbb{N}}$ of intervals such that
@@
where $|I_{n}|$ is the length of $I_{n}$. It is easy to see that every countable set has strong measure zero. Borel conjectured that a set is strong measure zero if and only if it is countable. The work of Sierpiński [Sierpiński1928] and Laver [16] proved that this is independent of ZFC. Strong measure zero is easily defined for subsets of Cantor space which we focus on in this paper. For a thorough background on strong measure zero sets see [2].
@@
In this work we investigate requiring the supermartingales themselves to be effective, leading instead to a stronger notion than strong measure zero sets in the classical setting. These supermartingale objects are similar to those used to define effective measure and dimension, but we will show in section 3 and that there are equivalent characterizations using the success rates of standard supermartingales.
@@
A natural effective version of Borel’s conjecture with time and space bounded resources along with at the computable level exists. This uses an effective countability definition by Lutz [20] which are roughly sets of sequences than can be uniformly enumerated within the resource bound. We prove that the effectively countable sets have effective strong measure zero at every level. For the time and space bounded setting it is proven that Borel’s conjecture holds. However, it does not hold at the computable level.
We investigate algorithmic strong measure zero occurring at the level of lower semicomputability. The main result is that the resulting class is equivalent to the well studied set of NCR reals defined by Reimann and Slaman [28] containing sequences that are not algorithmically random with respect to any continuous probability measure. We will discuss this connection more in section 5, but note for now that in the classical setting these are not equivalent. NCR has been proven to be countable [29]. We thus have the following implication diagram where $\mathrm{\Delta}_{R}$ represents a time or space bounded resource class and the dashed arrow implies independence from ZFC.
@@
We show that effective strong measure zero sets can be defined using effective coverings resulting in a characterization similar to Borel’s original definition (1.1). This results in another relatively simple characterization of NCR. While there are analagous classical results as well as in Higuchi and Kihara’s version, our setting requires use of different techniques. Using this result, we prove a correspondence principle giving sufficient conditions for computable, algorithmic, and classical versions of strong measure zero to agree on a set. This result is analogous to Hitchcock’s [13] correspondence principle for effective dimension. We use our result to show that there are sequences at arbitrarily high levels of the hyperarithmetic hierarchy with computable strong measure zero, mirroring a result for NCR.
Many of the characterizations of strong measure zero sets involve success with respect to a limit superior. We therefore investigate what happens if success is required in the limit inferior. For dimension, Lutz [24] showed a limit superior condition on martingales characterizes Hausdorff dimension while Althreya, Hitchcock, Lutz and Mayordomo [1] showed packing dimension can be characterized with a limit inferior. Requiring limit inferior success at all gauges leads to sets with strong packing dimension zero (SPDZ). In the classical setting, the existence of uncountable strong packing dimension zero sets is also independent of ZFC, see for example [34]. For time and space restrictions the sets are the same as strong measure zero. However at the computable level, the sets now correspond with a weak computably-countable notion that we define. It is also shown that the algorithmic version contains exactly the decidable sequences, giving the following implication chart.
@@ 2 Preliminaries
We work in the Cantor space $\mathbf{C}=\{0,1\}^{\infty}$ consisting of all infinite binary sequences. A string is a finite, binary string $w\in\{0,1\}^{*}$ which has length $|w|$. We write $\lambda$ for the empty string of length 0. For $0\leq i\leq j\leq|w|-1$ we write $w[i\dots j]$ for the string consisting of the $i^{\text{th}}$ through $j^{\text{th}}$ bits of $w$. Similarly for $i,j\in\mathbb{N}$ with $i\leq j$, we let $S[i\dots j]$ be the string consisting of the $i^{\text{th}}$ through $j^{\text{th}}$ bits of $S$.
@@
Fix an injective function $\langle\kern 5.0pt,\kern 5.0pt\rangle$ from $\{0,1\}^{*}\times\{0,1\}^{*}$ onto $\{0,1\}^{*}$ such that $\langle x,y\rangle$ and the projections $\langle x,y\rangle\to x,\langle x,y\rangle\to y$ are all computable in polynomial time.
@@
    $T=\bigcup_{n\in\mathbb{N}}\{f:\{0,1\}^{*}\to\{0,1\}^{*}\mid f\in\mathrm{DTIME}(g_{n})\}$.
Space bounded resource classes are defined analogously with DSPACE. In this paper we will let $\mathrm{\Gamma}_{R}$ be the set consisting of all time or space bounded resource classes that contain either P or PSPACE.
@@
$\mathrm{all}$ $=\{f:\{0,1\}^{*}\to\{0,1\}^{*}\}$
$\mathrm{comp}$ $=\{f\in\mathrm{all}\mid f\text{ is computable}\}$
@@
The result of a constructor $\delta$ is the sequence $R(\delta)\in\mathbf{C}$} such that
$\delta^{i}(\lambda)\sqsubseteq R(\delta)$ for all $i\in\mathbb{N}$.
@@
[…]
@@
Similarly, for $\mathrm{\Delta}\in\mathrm{\Gamma}_{RCA}$ and a discrete domain $D$ we say that $f:D\to[0,\infty)$ is $\mathrm{\Delta}$-computable if there is an function $g:\mathbb{N}\times D\to\mathbb{Q}\cap[0,\infty)\in\mathrm{\Delta}$ called a computation of $f$ such that $|g(r,x)-f(x)|\leq 2^{-r}$} for all $r\in\mathbb{N}$. In this paper $\mathrm{\Gamma}$ denotes $\mathrm{\Gamma}_{RCA}$ together with the lower semicomputable class. We will also refer to the lower semicomputable level as algorithmic.
@@
where $F^{\prime}(f)=g$ is a computation of $F(f)$. For $\mathrm{\Delta}\in\mathrm{\Gamma}$ equal to comp or lower semicomputable, we say that the functional $F$ is in $\mathrm{\Delta}$ if there is an oracle Turing machine $M$ such that for all $f:A\to B,r\in\mathbb{N}$ and $x\in D$, $M$ on input $(r,x)$ with oracle access to $f$ outputs $g(r,x)$ where $g\in\mathrm{\Delta}$ is a computation of $F(f)$ as defined above.
For time bounded $\mathrm{\Delta}$ we require that the time $M$ takes to output $g(x,r)$ is $h(|x|+r+\bar{g}(h(|x|+r))$ for some $h$ in $\mathrm{\Delta}$ where $\bar{g}(n)$ is the maximum length of the oracle $g$’s output on a string of length at most $n$. For example if $\mathrm{\Delta}$ is polynomial time then $M$ can query polynomially far out in the input length and use the maximum length of the query results as a parameter for the polynomial running time. Different notions of functional complexity exist, but all the results in the paper hold for every natural definition and most oracles will not impact the complexity in our use cases.
@@ 3.1 Strong Measure Zero
An odds function is any function $\{0,1\}^{*}\rightarrow[1,\infty). OO$ is said to be acceptable if $\prod_{w\sqsubset S}O(w)=\infty$ for all $S\in\mathbf{C}$.
@@
A set $X\subseteq\mathbf{C}$ has strong measure zero if and only if, for every acceptable odds $O:\{0,1\}^{*}\to[1,\infty)$, there is an $O$-supermartingale that succeeds on $X$.
@@
$\mu^{*}(X)=\inf\{\sum_{w\in A}\mu(w)\mid A\subseteq\{0,1\}^{*}\text{ and }X\subseteq\bigcup_{w\in A}C_{w}\}.$ (3.2)
@@
One can then show that there is an $O$-supermartingale that succeeds on a set $X$ if and only if $\mu_{O}^{*}(X)=0$. In their paper, they defined a set as having effective strong measure zero if for every computable odds function $O$ there exists an $O$-supermartingale that succeeds on the set. Here we instead use the following definition forcing effectivity on the $O$-supermartingales themselves.
@@
The fact that we limit odds functions to rationals in the interval $[1,2]$ may appear to be a weaker condition, but we will see that the definition is robust to this change and others later in this section. Note that all our odds functions therefore correspond to outer premeasures as in Lemma 3.3. From now on we will assume all odds functions are acceptable unless stated otherwise. We can also now see the connection between effective strong measure zero, measure zero, and dimension zero.
@@ 3.2 Strong Dimension Zero
$\limsup_{n\to\infty}\frac{d(S[0\dots n-1])}{2^{n}g(n)}=\infty$
@@
Suppose $X$ has $\mathrm{\Delta}$-strong measure zero with witnessing odds functional $F$. We will create a gauge functional $G$ as follows. Given a gauge function $g$, without loss of generality suppose $g(n)\leq 2g(n+1)$ for all $n$, otherwise $g$ can be transformed within $\mathrm{\Delta}$’s restrictions into a harder $g^{\prime}$ satisfying this with $g^{\prime}(n)\geq g(n)$ for all $n$. Now let $G$ use $F$ with odds function $O$ where $O(w)=\frac{g(|w|-1)}{g(|w|)}$ for all strings $w$ with $|w|>0$ and g(1)=1g(-1)=1. Then we have $\mu_{O}(w)=g(|w|)$ and
$G^{g}(w)=F^{O}(w)2^{|w|}g(|w|)$
@@
$\limsup_{n\to\infty}\frac{G^{g}(S[0\dots n])}{2^{n}g(n)}=\limsup_{n\to\infty}F^{O}(S[0\dots n])=\infty$
@@
$g(n)=\max\{\mu_{O}(w)\mid w\in\{0,1\}^{n}\}.$
Note that by compactness and the fact that $O$ is acceptable we have $lim_{n\to\infty}g(n)=0$. Therefore,
$F^{O}(w)=\frac{G^{g}(w)}{2^{|w|}\mu_{O}(w)}$
@@
$\limsup_{n\to\infty}F^{O}(S[0\dots n])\geq\limsup_{n\to\infty}\frac{G^{g}(S[0\dots n])}{2^{n}g(n)}=\infty$
@@
Let $D$ be any of the following sets or its intersection with $\mathbb{Q}$. Then, defining $\mathrm{\Delta}$ strong measure zero as the sets succeeding on all acceptable odds functions $O:\{0,1\}^{*}\to D$ results in equivalent definitions for every $\mathrm{\Delta}\in\mathrm{\Gamma}$.
@@ 3.3 Strong Predictability
$\liminf_{n\to\infty}\frac{\mathcal{L}^{\log}_{\pi}(S[0\dots n-1])}{h(n)}<1$
@@
$\iff$ $\liminf_{n\to\infty}n+\log g(n)-\log d(S[0\dots n])=$-$\infty\$
$\iff$ $\liminf_{n\to\infty}\mathcal{L}^{\log}_{\pi_{d}}(S[0\dots n])+\log g(n)=$-$\infty\$
$\iff$ $\liminf_{n\to\infty}\frac{\mathcal{L}^{\log}_{\pi_{d}}(S[0\dots n])}{-\log g(n)}<1\$
noting that the last equivalence is true since $\log g(n)$ is nonincreasing and goes to -$\infty$ while $\mathcal{L}^{\log}_{\pi_{d}}(S[0\dots n])$ is positive. ∎
@@
Let $\mathrm{\Delta}\in\mathrm{\Gamma}$. A set $X\subseteq\mathbf{C}$ has $\Delta$-Hausdorff measure zero with respect to a gauge $g$ if and only if there is a $\mathrm{\Delta}$-computable superpredictor that $\log(g)-$succeeds on $X$.